There are 6 men and 5 women. Of them, there will be 3 officers and no more than 2 men.
Answer by Boreal(15235)

2 men case: 6C2*5C1=15*5=75
1 man case: 6C1*5C2=6*10=60
0 men case: 5C3=10
145 ways.